The Brooks family and the Wong family each used their sprinklers last summer. The water output rate for the Brooks family's sprinkler was 40L per hour. The water output rate for the Wong family's sprinkler was 35L per hour. The families used their sprinklers for a combined total of 70 hours, resulting in a total water output of 2650L. How long was each sprinkler used?
Answer by ikleyn(52864)

.

Let x be the time for Brooks, and y be the time for Wong.


The system of equations is


     x  +    y = 70       (1)  (the combined time)

    40x + 35*y = 2650     (2)  (liters total


To solve it, multiply equation (1) by 35 (both sides).  Keep equation (2) as is


    35x + 35y = 70*35      (3)

    40x + 35y = 2650       (4)


Now subtract equation (3) from equation (4). You will get


    40x - 35 x = 2650-70*35

      5x       =  200

       x       = 200/5 = 40.


ANSWER.   Brooks family - 40 hours; Wong family  70-40 = 30 hours.

Solved.

----------------

If you want to see many other similar solved problems, look into the lesson
    - The Robinson family and the Sanders family each used their sprinklers last summer
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.